OpenStudy (anonymous):

anyone know how to do: -5.3= L(over)4 or L/4 ? plz help :/

OpenStudy (anonymous):

are you solving for L?

mathslover (mathslover):

\[\LARGE{-5.3=\frac{L}{4}}\] ? this is the question .. ?

OpenStudy (ganpat):

isolate L ... Put 4 on other side.. so that it will multiply over there.. -5.3 = L/4 So L = (4) * (-5.3)...

OpenStudy (anonymous):

thank u :)

mathslover (mathslover):

by the way .. \[\huge{\mathbb{Welcome}\textbf{To}\textbf{Open}\mathbb{Study}\textbf{@helloxox }}\]

OpenStudy (anonymous):

thanx :)

OpenStudy (anonymous):

this site is so helpful :)

OpenStudy (ganpat):

no one did such a warm welcome for me .. :(...

mathslover (mathslover):

nice to hear .. @helloxox For any help you can ask us .. and also for more information ... go to this link : http://openstudy.com/code-of-conduct sorry @Ganpat , i think when you enetered this site i was not an amby :)

OpenStudy (ganpat):

@mathslover : relax dude.. just kidding around.. I heard somewhere, its good to solve maths, with smile on ur face.. :D :-p

mathslover (mathslover):

..:-p "thanks for relaxation to me " .. well u heard right :)